# Insights into the Production of 44Ti and Nickel Isotopes in Core-Collapse Supernovae

### v1

@@ 4 Conclusion
Our 3D models produce 44Ti ranging from $10^{-6}M_{\odot}$ to $2\times 10^{-4}M_{\odot}$. The high 44Ti/56Ni ratio observed in Cas A can easily be achieved by many of our models. Therefore, the “44Ti-problem” (The et al., 2006) is no longer an issue in self-consistent long-term, 3D, initially non-rotating CCSN models.
@@ Acknowledgments
This paper uses previously published data in https://doi.org/10.5281/zenodo.10498614. This online repository includes all the tracer trajectories (masses, positions, temperatures, electron fractions, and local neutrino spectra), together with the ejecta abundances of each model at the end of the 3D simulations. The progenitor models from Sukhbold et al. (2016) and Sukhbold et al. (2018) are available as well. The 17, 18, 18.5, 20, 25, and 60 $M_{\odot}$ models have been evolved to later times in 3D since the publication of the above repository. However, we can not replace the files in the repository, as this would significantly exceed to size limit of the repository. Therefore, we encourage readers to start with the online repository, and if necessary, to make direct requests to the authors for the rest of data.

### v2

@@ 4 Conclusion
Our 3D models produce 44Ti ranging from $10^{-6}M_{\odot}$ to $2\times 10^{-4}M_{\odot}$. The high 44Ti/56Ni ratio observed in Cas A can be achieved by some of our models, though the data point is still on the high side. Given the fact that Cas A may still contain some unshocked Fe-rich matter, the actual 44Ti/56Ni ratio might be a bit lower, which would then be even more consistent with our models. Therefore, the “44Ti-problem” (The et al., 2006) is no longer an issue in self-consistent long-term, 3D, initially non-rotating CCSN models.
@@ Acknowledgments
This paper uses previously published data, which can be accessed viahttps://doi.org/10.5281/zenodo.10498614 (catalog DOI: 10.5281/zenodo.10498614) (Wang, 2024). This online repository includes all the tracer trajectories (masses, positions, temperatures, electron fractions, and local neutrino spectra), together with the ejecta abundances of each model at the end of the 3D simulations. The progenitor models from Sukhbold et al. (2016) and Sukhbold et al. (2018) are available as well. The 17, 18, 18.5, 20, 25, and 60 $M_{\odot}$ models have been evolved to later times in 3D since the publication of the above repository. However, we can not replace the files in the repository, as this would significantly exceed to size limit of the repository. Therefore, we encourage readers to start with the online repository, and if necessary, to make direct requests to the authors for the rest of data.